# Bayesian-LoRA: Probabilistic Low-Rank Adaptation of Large Language Models

### v1

@@ 2 Related Work
Bayesian methods quantify uncertainty via BNNs (Izmailov et al., 2021; Lin et al., 2025a; Kweon et al., 2025), Laplace approximations (Yang et al., 2024), and Gaussian processes (Ranković & Schwaller, 2025), but incur high cost at LLM scale. Full-model methods (including BLoB (Wang et al., 2024b)) require approximate inference over entire parameter spaces (Graves, 2011; Blundell et al., 2015); Laplace methods depend on Hessians whose cost grows with model size (Daxberger et al., 2021; Ritter et al., 2018); GPs scale cubically in data (Seeger, 2004; Quinonero-Candela & Rasmussen, 2005); and ensembles multiply inference cost (Lakshminarayanan et al., 2017). We instead model uncertainty within the low-rank subspace of LoRA, combining sparse GP with normalizing flows to induce $\Delta W$ with low overhead.
@@ 3.2 Variational Sparse Inducing Weight Model
Rather than placing a distribution directly on $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, we introduce a compact inducing matrix $U\in\mathbb{R}^{r\times c}$ with $r\ll d_{\text{out}}$, $c\ll d_{\text{in}}$ that controls the distribution over $W$ (Yang et al., 2024; Snelson & Ghahramani, 2005). As in Sparse GPs, the low-dimensional $U$ acts as a sufficient statistic for the posterior over $W$, keeping inference tractable.
@@ A.3 ELBO Derivation
Starting from $\log p(\mathcal{D})=\log\!\int p(U,W,\mathcal{D})\,\mathrm{d}U\,\mathrm{d}W$ and inserting $q(U,W)$:

### v2

@@ 2 Related Work
Bayesian methods quantify uncertainty by BNNs (Izmailov et al., 2021; Lin et al., 2025a; Kweon et al., 2025), Laplace approximations (Yang et al., 2024), and Gaussian processes (Ranković & Schwaller, 2025), but incur high cost at LLM scale. Full-model methods (including BLoB (Wang et al., 2024b)) require approximate inference over entire parameter spaces (Graves, 2011; Blundell et al., 2015); Laplace methods depend on Hessians whose cost grows with model size (Daxberger et al., 2021; Ritter et al., 2018); GPs scale cubically in data (Seeger, 2004; Quinonero-Candela & Rasmussen, 2005; Ritter et al., 2021); and ensembles multiply inference cost (Lakshminarayanan et al., 2017). FFG-U (Ritter et al., 2021) proposed inducing weights by Kronecker-factorized decomposition of SGP weights that introduces a low-dimensional inducing weight matrix U. We therefore inspired model uncertainty within the low-rank subspace of LoRA, combining sparse GP with normalizing flows to induce $\Delta W$ with low overhead.
@@ 3.2 Variational Sparse Inducing Weight Model
Rather than placing a distribution directly on $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, we introduce a compact inducing matrix $U\in\mathbb{R}^{r\times c}$ with $r\ll d_{\text{out}}$, $c\ll d_{\text{in}}$ that controls the distribution over $W$ (Ritter et al., 2021; Yang et al., 2024; Snelson & Ghahramani, 2005). As in Sparse GPs, the low-dimensional $U$ acts as a sufficient statistic for the posterior over $W$, keeping inference tractable.
@@ A.3 ELBO Derivation
Starting from $\log p(\mathcal{D})=\log\!\int p(U,W,\mathcal{D})\,\mathrm{d}U\,\mathrm{d}W$ and inserting $q(U,W)$ (Ritter et al., 2021):